# Long-range Meta-path Search on Large-scale Heterogeneous Graphs

### v4

@@ 5.1 Progressive Sampling Search
Let =$\mathbb{P}=\{P_{1},\cdots,P_{k},\cdots,P_{K}\}$ be the initial search space with all $K$ target-node-related meta-paths. Let $\bm{X}^{c_{i}}$ be the raw feature matrix of all nodes belonging to type $c_{i}$, and $\bm{\hat{A}}_{c_{i},c_{i}+1}$ be the row-normalized adjacency matrix between node type $c_{i}$ and $c_{i+1}$. The neighbor aggregation block follows SeHGNN (Yang et al., 2023), which pre-process the heterogeneous graph into regular-shaped tensors for target nodes. Specifically, it uses the multiplication of adjacency matrices to calculate the final contribution weight of each metapath-based neighbor to targets. As shown in Figure 2, the neighbor aggregation process of $l$-hop meta-path $P_{k}=c_{0}c_{1}c_{2}\ldots c_{l}$ is:
@@
To automatically discover meaningful meta-paths for various datasets or tasks without prior, our search space contains all target-node-related meta-paths, severely challenging the efficiency and effectiveness of searching. To overcome the efficiency challenge, LMSPS utilizes a progressive sampling algorithm to sample a part of meta-paths in each iteration and progressively shrink the search space.
Specifically, LMSPS assigns one architecture parameter to each meta-path. Let $\bm{\alpha}=\{\alpha_{1},\cdots,\alpha_{k},\cdots,\alpha_{K}\}\in\mathbb{R}^{K}$ be corresponding architecture parameters of $K$ meta-paths. We use a Gumbel-softmax (Maddison et al., 2016; Dong & Yang, 2019) over architecture parameters $\alpha_{k}$ to calculate the strength of different meta-paths:
@@
where $q_{k}$ is the path strength, representing the relative importance of meta path $P_{k}$. $u_{k}=-\log\left(-\log\left(U\right)\right)$ where $U\sim\operatorname{Uniform}(0,1)$, and $\tau$ is the temperature controlling the continuous relaxation’s extent.
The progressive sampling algorithm uses the path strength to progressively narrow the search space from $K$ to $V$ to exclude useless meta-paths. Generally, $K\gg V$ under a large maximum hop. Let $\widetilde{q}_{C}$ be the $C$-th largest path strength of =$\mathbb{Q}=\{q_{1},\cdots,q_{k},\cdots,q_{K}\}$. During the search stage, the search space retains all meta-paths no less than $\widetilde{q}_{C}$. The dynamic search space can be formulated as:
$\mathbb{S}_{C}$ =$=\{k|q_{k}\geq\widetilde{q}_{C},\forall 1\leq k\leq K\},$ (3)
$\text{where}\,\,C$ $=\lceil\lambda(K-V)\rceil+V.$ (4)
@@
As the search stage aims to determine top-$M$ meta-paths, we sample $M$ meta-paths from dynamic search space in each iteration. In each iteration, only parameters on the $M$ activated meta-paths are updated, and the rest remain unchanged. Therefore, the search cost is relevant to $M$ instead of $K$. The forward propagation can be expressed as:
$\bm{Z}$ $={\rm MLP}\Big{(}\sum_{k\in\mathbb{S}}q_{k}^{\prime}\cdot{\rm MLP}_{k}(\bm{X}_{k})\Big{)},$ (5)
$\text{where}\,\,\mathbb{S}$ $={\rm UniformSample}(\mathbb{S}_{C},M).$ (6)
@@
The parameter update in the super-net involves a bilevel optimization problem (Anandalingam & Friesz, 1992; Colson et al., 2007; Xue et al., 2021).
$\min_{\bm{\alpha}}\,\,$ $\mathcal{L}_{val}(\mathbb{S},\bm{\omega}^{*}(\bm{\alpha}),\bm{\alpha}),$ (7)
s.t. $\bm{\omega}^{*}(\bm{\alpha})=\mathrm{argmin}_{\bm{\omega}}\,\mathcal{L}_{train}(\mathbb{S},\bm{\omega},\bm{\alpha}).$ (8)
Here $\mathcal{L}_{train}$ and $\mathcal{L}_{val}$ denote the training and validation loss, respectively. $\bm{\alpha}$ is the architecture parameters calculating the path strength. $\bm{\omega}$ is the network weights in MLP. Following the NAS-related works in the computer vision field (Liu et al., 2019b; Xie et al., 2019; Yao et al., 2020), we address this issue by first-order approximation. Specifically, we alternatively freeze architecture parameters $\bm{\alpha}$ when training $\bm{\omega}$ on the training set and freeze $\bm{\omega}$ when training $\bm{\alpha}$ on the validation set.
The progressive sampling strategy can contribute to a more compact search space specifically driven by the current HIN and task, leading to a more effective meta-path discovery. Additionally, it can overcome the deep coupling issue (Guo et al., 2020) because of the randomness in each iteration.
HAN $O(N(rd)^{l}F^{2})$ $O(N(rd)^{l}F)$ $O(Nr^{l}F^{2})$ $O(N(rd)^{l}F^{2})$
Simple-HGN $O(N(rd)^{l}F^{2})$ $O(N(rd)^{l}F)$ $O(N(rd)^{l}F^{2})$
${}^{\dagger}$ $O(NrdlF^{2})$ $O(NrdlF)$ $O(NrdlF^{2})$
SeHGNN $O(Nr^{l}F^{2})$ $O(Nr^{l}F^{2}+Nr^{2l}F)$ $O(N(r^{l}F^{2}+r^{2l}F))$
LMSPS-search $O(NMF^{2})$ $O(NMF^{2})$ $O(NMF^{2})$
LMSPS-train $O(NMF^{2})$ $O(NMF^{2})$ $O(NMF^{2})$
@@ 5.2 Sampling Evaluation
After the completion of progressive sampling search, the search space is narrowed from $K$ to $V$. Traditional methods in the computer vision field directly derive the final architecture based on the architecture parameters (Liu et al., 2019b; Xie et al., 2019; Yao et al., 2020). However, a recent research (Wang et al., 2021) has verified that architecture parameters are not effective enough to indicate the strength of the candidates. In addition, because different meta-paths could be noisy or redundant to each other, top-$M$ meta-paths are not necessarily the optimal solution when their importance is calculated independently. Based on this consideration, we propose a novel sampling evaluation strategy. Specifically, using path strength as the probability, we sample $M$ meta-paths from the compact search space $\mathbb{S}_{V}$ to evaluate their overall performance. The sampling evaluation is repeated $200$ times to filter out the meta-path set with the lowest validation loss. This stage is not time-consuming because the evaluation does not involve weight training. This sampling process can be represented as:
@@
Here, DiscreteSample$(\mathbb{S}_{V},M,\bar{\mathbb{Q}})$ indicates a set of $M$ elements chosen from the set $\mathbb{S}_{V}$ without replacement via discrete probability distribution $\bar{\mathbb{Q}}$. $\bar{\mathbb{Q}}$ is the set of relative path strength calculated by architecture parameters of the $V$ meta-paths based on Equation 2. The overall search algorithm and more discussion are shown in Appendix B.
@@
5.3 Time Complexity Analysis
Following the convention (Chiang et al., 2019; Yang et al., 2023), we compare the time complexity of LMSPS with HAN (Wang et al., 2019), Simple-HGN (Lv et al., 2021), and SeHGNN (Yang et al., 2023) under mini-batch training with the total $N$ target nodes. All methods employ $l$-hop neighborhood. For simplicity, we assume that the number of features is fixed to $F$ for all layers. The average degree of each node is $rd$, where $r$ is the number of edge types and $d$ is the number of edges connected to the node for each edge type. The complexity analysis is summarized in Table 1. Because HAN and Simple-HGN require neighbor aggregation during training, and the number of neighbors grows exponentially with hops. So, they have neighbor aggregation costs of $O((rd)^{l})$. SeHGNN employs a pre-processing step to avoid the training cost of neighbor aggregation. However, the exponential meta-paths cause SeHGNN to suffer from $O(r^{l})$ costs in semantic aggregation. Unlike the above methods, LMSPS samples $M$ meta-paths in each iteration of the search stage and employs $M$ effective meta-paths in the training stage to avoid exponential costs. Generally, we have $O((rd)^{l})\gg O(r^{l})=O(K)\gg O(M)$ when maximum hop $l$ is large. The time complexity of LMSPS is a constant when $N$ and $F$ are determined, which is the key point for utilizing long-range meta-paths.
@@ 6 Experiments and Analysis
This section evaluates the benefits of our method against state-of-the-art models on eight heterogeneous datasets. We aim to answer the following questions: Q1. How does LMSPS perform compared with state-of-the-art baselines? Q2. Does LMSPS perform better on sparser heterogeneous graphs? Q3. Can LMSPS overcome the over-smoothing and noise issues? Q4. Are the searched meta-paths effective on other models? More ablation studies and hyper-parameter studies are provided in Appendix D.6 and D.7.
[…]
Method Test accuracy
Random $35.14\pm 3.78$
MLP (Hu et al., 2020a) $26.92\pm 0.26$
GraphSAGE (Hamilton et al., 2017) $46.78\pm 0.67$
RGCN (Schlichtkrull et al., 2018) $47.37\pm 0.48$
HGT (Hu et al., 2020b) $49.27\pm 0.61$
NARS (Yu et al., 2020) $50.88\pm 0.12$
SeHGNN (Yang et al., 2023) $51.79\pm 0.24$
[…]
SAGN (Sun et al., 2021)+label $51.17\pm 0.32$
GAMLP (Zhang et al., 2022)+label $51.63\pm 0.22$
SeHGNN+label $53.99\pm 0.18$
LMSPS +label $\textbf{55.10}\pm\textbf{ 0.11}$
SeHGNN+label+ms $56.71\pm 0.14$
LMSPS +label+ms $\textbf{57.84}\pm\textbf{ 0.22}$
Method DBLP IMDB ACM
RSHN (Zhu et al., 2019) $93.81\pm 0.55$ $64.22\pm 1.03$ $90.32\pm 1.54$
HetSANN (Hong et al., 2020) $80.56\pm 1.50$ $57.68\pm 0.44$ $89.91\pm 0.37$
GTN (Yun et al., 2019) $93.97\pm 0.54$ $65.14\pm 0.45$ $91.20\pm 0.71$
HGT (Hu et al., 2020b) $93.49\pm 0.25$ $67.20\pm 0.57$ $91.00\pm 0.76$
Simple-HGN (Lv et al., 2021) $94.46\pm 0.22$ $67.36\pm 0.57$ $93.35\pm 0.45$
HINormer (Mao et al., 2023) $94.94\pm 0.21$ $67.83\pm 0.34$-
RGCN (Schlichtkrull et al., 2018) $92.07\pm 0.50$ $62.05\pm 0.15$ $91.41\pm 0.75$
HetGNN (Zhang et al., 2019) $92.33\pm 0.41$ $51.16\pm 0.65$ $86.05\pm 0.25$
HAN (Wang et al., 2019) $92.05\pm 0.62$ $64.63\pm 0.58$ $90.79\pm 0.43$
MAGNN (Fu et al., 2020) $93.76\pm 0.45$ $64.67\pm 1.67$ $90.77\pm 0.65$
SeHGNN (Yang et al., 2023) $95.24\pm 0.13$ $68.21\pm 0.32$ $93.87\pm 0.50$
DiffMG (Ding et al., 2021) $94.20\pm 0.36$ $59.75\pm 1.23$ $88.07\pm 3.04$
PMMM (Li et al., 2023) $95.14\pm 0.22$ $67.58\pm 0.22$ $93.71\pm 0.17$
LMSPS $\textbf{95.66}\pm\textbf{0.20}$ $\textbf{68.70}\pm\textbf{0.26}$ $\ \textbf{94.69}\pm\textbf{0.36}$
[…]

### v6

@@ 5.1 Progressive Sampling Search
Let =$\mathbb{P}=\{P_{1},\cdots,P_{k},\cdots,P_{K}\}$ be the initial search space with all the $K$ target-node-related meta-paths, $\bm{X}^{c_{i}}$ be the raw feature matrix of all nodes belonging to type $c_{i}$, and $\bm{\hat{A}}_{c_{i},c_{i}+1}$ be the row-normalized adjacency matrix between node type $c_{i}$ and $c_{i+1}$. The neighbor aggregation block follows SeHGNN [51], which pre-processes the heterogeneous graph into regular-shaped tensors for target nodes. Specifically, it uses the multiplication of adjacency matrices to calculate the final contribution weight of each metapath-based neighbor to targets. As shown in Figure 2, the neighbor aggregation process of $l$-hop meta-path $P_{k}=c_{0}c_{1}c_{2}\ldots c_{l}$ is:
@@
To automatically discover meaningful meta-paths for various datasets or tasks without prior, our search space contains all target-node-related meta-paths, severely challenging the efficiency and effectiveness of the search. To address the efficiency challenge, LMSPS utilizes a progressive sampling algorithm to sample meta-paths in each iteration and progressively shrink the search space.
Specifically, LMSPS assigns one architecture parameter to each meta-path. Let $\bm{\alpha}=\{\alpha_{1},\cdots,\alpha_{k},\cdots,\alpha_{K}\}\in\mathbb{R}^{K}$ be the architecture parameters of $K$ meta-paths. We use a Gumbel-softmax [33, 8] over architecture parameters $\alpha_{k}$ to calculate the strength of different meta-paths:
@@
where $q_{k}$ is the path strength, representing the relative importance of meta path $P_{k}$. $u_{k}=-\log\left(-\log\left(U\right)\right)$ where $U\sim\operatorname{Uniform}(0,1)$, and $\tau$ is temperature controlling relaxation’s extent.
The progressive sampling algorithm uses the path strength to progressively narrow the search space from $K$ to $V$ to exclude useless meta-paths. Generally, $K\gg V$ under a large maximum hop. Let $\widetilde{q}_{C}$ be the $C$-th largest path strength of =$\mathbb{Q}=\{q_{1},\cdots,q_{k},\cdots,q_{K}\}$. During the search stage, the search space retains all meta-paths no less than $\widetilde{q}_{C}$. The dynamic search space can be formulated as follows:
$\mathbb{S}_{C}=\{k|q_{k}\geq\widetilde{q}_{C},\forall 1\leq k\leq K\}\quad\text{where}\,\,\,C=\lceil\lambda(K-V)\rceil+V.$ (3)
@@
As the search stage aims to determine top-$M$ meta-paths, we sample $M$ meta-paths from dynamic search space in each iteration. In each iteration, only parameters on the $M$ activated meta-paths are updated, while others remain unchanged. Therefore, the search cost is relevant to $M$ instead of $K$. The forward propagation can be expressed as:
$\bm{Z}={\rm MLP}\Big{(}\sum_{k\in\mathbb{S}}q_{k}^{\prime}\cdot{\rm MLP}_{k}(\bm{X}_{k})\Big{)}\quad\text{where}\,\,\,\mathbb{S}={\rm UniformSample}(\mathbb{S}_{C},M).$ (4)
@@
The parameter update in the super-net involves a bilevel optimization problem [2, 6, 50].
$\min_{\bm{\alpha}}\,\,\mathcal{L}_{val}(\mathbb{S},\bm{\omega}^{*}(\bm{\alpha}),\bm{\alpha})\quad\text{s.t.}\,\,\,\bm{\omega}^{*}(\bm{\alpha})=\mathrm{argmin}_{\bm{\omega}}\,\mathcal{L}_{train}(\mathbb{S},\bm{\omega},\bm{\alpha}).$ (5)
Here $\mathcal{L}_{train}$ and $\mathcal{L}_{val}$ denote the training and validation loss, respectively. $\bm{\alpha}$ is the architecture parameters calculating the path strength. $\bm{\omega}$ is the network weights in MLP. Following the NAS-related works in the computer vision field [31, 49, 52], we address this issue by first-order approximation. Specifically, we alternatively freeze architecture parameters $\bm{\alpha}$ when training $\bm{\omega}$ on the training set and freeze $\bm{\omega}$ when training $\bm{\alpha}$ on the validation set.
The progressive sampling strategy can contribute to a more compact search space specifically driven by the current HIN and task, leading to a more effective meta-path discovery. Additionally, it can overcome the deep coupling issue [12] because of the randomness in each iteration.
@@ 5.2 Sampling Evaluation
After the completion of the progressive sampling search, the search space is narrowed from $K$ to $V$. Traditional methods in the computer vision field directly derive the final architecture based on the architecture parameters [31, 49, 52]. However, as different meta-paths could be noisy or redundant to each other, top-$M$ meta-paths are not necessarily the optimal solution when their importance is calculated independently. Based on this consideration, we specially designed a novel sampling evaluation strategy for meta-path search. Specifically, using path strength at the end of progressive sampling as the probability, we sample $M$ meta-paths from the compact search space $\mathbb{S}_{V}$ to evaluate their overall performance. The sampling evaluation is repeated $200$ times to filter out the meta-path set with the lowest validation loss. So, we can select the best meta-path set instead of independent top-$M$ meta-paths, which is more reasonable. This stage is not time-consuming because the evaluation does not involve weight training. This sampling process can be represented as:
@@
Here, DiscreteSample$(\mathbb{S}_{V},M,\bar{\mathbb{Q}})$ indicates a set of $M$ elements chosen from set $\mathbb{S}_{V}$ without replacement via discrete probability distribution $\bar{\mathbb{Q}}$. $\bar{\mathbb{Q}}$ is the set of relative path strength calculated by architecture parameters of the $V$ meta-paths based on Equation 2. The overall search algorithm and more discussion are shown in Appendix C.
@@
5.3 Discussion on Differences with Prior Works
As discussed in Section 3, there have been some attempts to find meta-paths automatedly [14, 55, 59]. Three aspects highlight the differences between LMSPS and existing methods. 1) Large-scale dataset: LMSPS is the first HGNN that makes it possible to achieve automated meta-path selection for large-scale heterogeneous graph node property prediction. 2) Long-range dependency: LMSPS is the first HGNN to utilize long-range dependency in large-scale heterogeneous graphs. To achieve the above two goals, LMSPS has addressed two key challenges: (a) alleviating costs while striving to effectively utilize information in exponentially increased receptive fields and (b) overcoming the well-known over-smoothing issue. 3) High generalization: Based on Table 5, the searched meta-paths of LMSPS can be generalized to other HGNNs to boost their performance, which has not been achieved by existing works. To accomplish this objective, LMSPS uses an MLP-based architecture instead of a transformer-based one for meta-path search because the former involves fewer inductive biases [43, 47], i.e., human interventions.
@@ 6 Experiments and Analysis
This section evaluates the benefits of our method against state-of-the-art models on nine heterogeneous datasets. We aim to answer the following questions: Q1. How does LMSPS perform compared with state-of-the-art baselines? Q2. Can LMSPS overcome the over-smoothing and noise issues? Q3. Are the search algorithm and searched meta-paths effective? Q4. Does LMSPS perform better on sparser heterogeneous graphs?